mass of one atom of X is 6.642 × 10^-23 g.What is its atomic mass?

Grade:11

2 Answers

Vikas TU
6 years ago
As we know that 1 mole of any atom= 6.022*10^23 atoms
          Or we can say that 6.022*10^23 atoms have mass=1 mol=atomic mass of
          that element.
          Now mass of atom= 6.642 *10^-23g
          Mass of 6.022*10^23 atoms = (6.022*10^23)*( 6.642 *10^-23g)
                                                              = 39.998124= 40u
Vikas TU
6 years ago
From 1 mole of any atom= 6.022*10^23 particles 
Or, on the other hand we can state that 6.022*10^23 iotas have mass=1 mol=atomic mass of 
that component. 
Presently mass of atom= 6.642 *10^-23g 
Mass of 6.022*10^23 molecules = (6.022*10^23)*( 6.642 *10^-23g) 
= 39.998124= 40u